Thank you so much for your help !!!!
There are 850 dougles fir and ponderosa pine trees in a section of forest bought by Sawz logging co. The company paid an average of 300.00 for each douglas fir and 225.00 for each ponderosa pine. If the company paid 217,500.00 for the trees how many of each kind did the company buy?

Answer by CharStar(110)

This is a mixture problems when you look at it clearly.
$300.00 For Douglas fir - $300x
$225.00 for Ponderosa pine = $225(850-x)
The equation that I get is $300x-225(850-x)=$217,500
$300-191,250-225x=$217,500
75x-191,250=$217,500
75x=$26,250
x=350
To Check your answer
350 Douglas fir is - $300.00(350) = $105,000
500 Ponderosa Pin is - $225(500) = $112,500
Total is $217,500
I hope that helps you
CharStar